OpenStudy (hellokitty17):

Simplify (4x - 6) + (5x + 1). a. 9x + 5 b.9x - 5 c.x - 5 d.-x - 5

Nnesha (nnesha):

there is plus sign so you can remove the parentheses and then combine like terms add coefficient of same variable and number without any variable combine them

Nnesha (nnesha):

for example 3x+x 3x and x are like terms so you combine them 3x+x =4x

OpenStudy (kendricklamar2014):

1. Simplify the equation:\[4x−6+5x+1 = 4x+−6+5x+1\] 2. Combine like terms: \[4x+−6+5x+1 = (4x+5x)+(−6+1)\] 3. Combine like terms: \[(4x+5x)+(−6+1) = 9x+−5\] Answer:\[9x−5\]

oregonduck (oregonduck):

It is B

OpenStudy (kendricklamar2014):

Yes, the answer is B

Nnesha (nnesha):

.

OpenStudy (kendricklamar2014):

,